I'm sorry! Can you help me this problems. I have been study about Complex Numbers i just studying at hight scholl .But i can't find this problem so i want you help me in this 2+3i, 4+3i,in form r(cosQ^+isinQ^).Thank you.
Answer by stanbon(75887)

2+3i, 4+3i,in form r(cosQ^+isinQ)
------------
Forms:
a+bi can be converted using the following:
r= sqrt(a^2+b2)
Q = tan^-1(b/a)
-----------------------------
Your Problems
2+3i converts as follows:
r = sqrt(2^2+3^2) = sqrt(13)
Q = tan^1(3/2) = 56.3 degrees
2+3i = (sqrt(13))(cos(56.3)+isin(56.3))
--------------------
4+3i converts as follows:
r = sqrt(4^2+3^2)=sqrt(25) = 5
Q = tan^-1(3/4) = 36.87 degrees
4+3i = 5(cos(36.87)+isin(36.87))
====================
Cheers,
Stan H.